OpenStudy (anonymous):

a rectangle has a volume of 3 cubic feet.the rectangle is 24 inches long and 18 inches wide.what is the height in inches of the rectangle

OpenStudy (anonymous):

What is the formula for the volume of a rectangle?

OpenStudy (anonymous):

One thing that I should point out is that 3 cubic feet= \[1ft^{3}*\frac{12in}{1ft}\frac{12in}{1ft}\frac{12in}{1ft}=1728in\]

OpenStudy (anonymous):

Do you know what to do?

OpenStudy (anonymous):

i know i have to find the height bt how do i find the height

OpenStudy (anonymous):

You remember the formula for the volume of a rectangular prism, right?

OpenStudy (anonymous):

length\(*\)width\(*\)height?

OpenStudy (anonymous):

yes bt can you answer the question that i put

OpenStudy (anonymous):

Yes, but I'm not just going to tell you the answer.

OpenStudy (anonymous):

What you need to do is realize that length*width*height=volume. So in this case, 24*18*something=1728. What does "something" have to be?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

why does it have to equal 1728

OpenStudy (anonymous):

Sorry, not 1728, 5184. 1 cubic foot is 1728 cubic inches, so 3 cubic feet is 5184 cubic inches.

OpenStudy (anonymous):

so then what do i have to do now

OpenStudy (anonymous):

Just replace 1728 by 5184 and figure out what you need instead of 4.

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

so the answer is 12

OpenStudy (anonymous):

can you help me with another question

OpenStudy (anonymous):

Yes, the answer is 12.

OpenStudy (anonymous):

Sure, what's the question?

OpenStudy (anonymous):

what is the volume of a cone shaped party hat with a base whose circumference is25.12 inches and height is 8 inches

OpenStudy (anonymous):

oh and i have many math questions to ask you if you dont mind

OpenStudy (anonymous):

hellooooooooooooooooo

OpenStudy (anonymous):

hello are you there